# Infinite-Horizon Optimal Wireless Control Over Shared State-Dependent Fading Channels for IIoT Systems

### v1

@@ 3.2 Lyapunov-like performance analysis of WCS
Given $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$, MAS (4) is constrained $\mathcal{M}$-stabilizable from initial state $\alpha_{0}$, if there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$ such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha}$ holds for any $k<K_{0}$, while $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{M}$ holds for any $k\geq K_{0}$.
@@
A subset $\mathcal{S}\subseteq\mathcal{C}_{\alpha}$ is called a constrained control invariant subset (CCIS) of MAS (4), if for any $\alpha(0)=\alpha_{0}\in\mathcal{S}$, there exists an admissible input sequence ${\bf{u}}$ such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{S}$ holds for any $k\in\mathbb{N}$.
Obviously, the union of any two CCISs of MAS (4) is another CCIS. The union of all CCISs contained in $\mathcal{M}$ is called its largest constrained control invariant set, denoted by $I(\mathcal{M})$.
@@
The Lyapunov-like performance (6) is guaranteed for WCS (3) with the state-dependent fading channel (5) and safety constraints $\mathcal{C}_{\alpha}$, $\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}}))$, if and only if MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, where $\Omega(s)=\{\alpha\in\mathcal{C}_{\alpha}:\bar{\lambda}_{i}(\alpha)\geq s_{i},i=1,\cdots,q\}$, $s=[s_{1}~{}\cdots~{}s_{q}]$ with
$s_{i}=\sup\limits_{y\in\mathbb{R}^{n_{i}\times 1},y\neq 0}\frac{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})y}{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})y}.$ (11)
@@
For each subsystem $i$, since $\mathbb{E}[\xi_{i}(k)]=\mathbf{0}$ and $\lambda_{i}(k)$ is independent of $x_{i}(k)$, by (3), it holds
$\mathbb{E}[V_{i}(x_{i}(k+1))|x_{i}(k),\alpha(k;\alpha_{0},{\bf{u}})]$
$=$ $\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}x_{i}^{\top}(k)A_{c,i}^{\top}Q_{i}A_{c,i}x_{i}(k)$
+$+\mathbb{P}\{\lambda_{i}(k)=0|\alpha(k;\alpha_{0},{\bf{u}})\}x_{i}^{\top}(k)A_{o,i}^{\top}Q_{i}A_{o,i}x_{i}(k)$
@@
Then, condition (6) subject to the multiloop WCS dynamics (3) is equivalent to the condition that for all $x_{i}(k)\neq\mathbf{0}$, it holds
$\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$
$\geq$ $\frac{x_{i}^{\top}(k)(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})x_{i}(k)}{x_{i}^{\top}(k)(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})x_{i}(k)},$ (12)
which is equivalent to $\bar{\lambda}_{i}(\alpha(k;\alpha_{0},{\bf{u}}))\geq s_{i}$ with $s_{i}$ given in (11), that is, MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$.
Next, we prove that MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$ if and only if it is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. We just need to prove the necessity. This is proved by a contradiction argument. Since MAS (4) is constrained $\Omega(s)$-stabilizable from $\alpha_{0}$, we can obtain a feasible admissible input sequence $\bar{{\bf{u}}}$ and a positive integer $K_{0}$, such that for any integer $k<K_{0}$, it holds $\alpha(k;\alpha_{0},\bar{{\bf{u}}})\in\mathcal{C}_{\alpha}$; for any integer $k\geq K_{0}$, it holds $\alpha(k;\alpha_{0},\bar{{\bf{u}}})\in\Omega(s)$.
Assuming that MAS (4) is not constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then there exists an integer $k_{0}\geq K_{0}$ satisfying $\alpha(k_{0};\alpha_{0},\bar{{\bf{u}}})\notin I(\Omega(s))$. Since $I(\Omega(s))$ is a largest CCIS, for any $\bar{\alpha}\in\Omega(s)\setminus I(\Omega(s))$ and any admissible input sequence ${\bf{u}}$, there exists an integer $k(\bar{\alpha},{\bf{u}})$ such that $\alpha(k(\bar{\alpha},{\bf{u}});\bar{\alpha},{\bf{u}})\notin\Omega(s)$. Then, it holds
$\alpha(k_{0}+k(\alpha(k_{0};\alpha_{0},\bar{{\bf{u}}}),\tilde{{\bf{u}}});\alpha_{0},\bar{{\bf{u}}})\notin\Omega(s),~{}\tilde{u}(k)=\bar{u}(k+k_{0}),$
which is a contradiction. Thus, MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. $\Box$
By (12), $s_{i}$ represents the lower bound of the probability of successful transmission for system $i$ that ensures the desired Lyapunov-like performance given in Definition 1 under the WCS and MAS coupling. In addition, $\Omega(s)$ consists of all constrained states of the MAS satisfying the performance requirements constraint (6) of WCS (3). Thus, for the optimal control problem (8), all constraints can be guaranteed if MAS (4) is constrained $I(\Omega(s))$-stabilizable, that is, problem (8) is equivalently converted into the following infinite-horizon optimal constrained set stabilization problem of MAS (4):
$\min_{{\bf{u}}}J_{{\bf{u}}}(\alpha_{0})$
@@
$~{}~{}~{}~{}~{}~{}~{}\,\exists\ K_{0}>0,\alpha(k;\alpha_{0},{\bf{u}})\in I(\Omega(s)),\forall\ k\geq K_{0},$
$~{}~{}~{}~{}~{}~{}~{}\,\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha},\forall\ k<K_{0},$
$~{}~{}~{}~{}~{}~{}~{}\,u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})),\forall\ k\in\mathbb{N}.$ (13)
4 Feasibility analysis
In this section, we first reformulate the MAS in a tractable form, and then present a criterion on the constrained set stabilization of MAS (4). Finally, the feasibility of optimal control problem (13) is analyzed based on the graphical method.
4.1 Algebraic reformulation of the MAS
Due to the highly nonlinear finite field model, it is difficult to apply classical nonlinear system theory to investigate (4). Recently, the semi-tensor product approach has emerged as an elegant tool to deal with such issues [chengqi2011, cheng2011]. The basic idea is to embed the nonlinear finite filed dynamics in (4) into a higher dimensional space where its evolution is linear. In this part, by virtue of the semi-tensor product, we convert MAS (4) into a bilinear equivalent algebraic form, which facilitates the sequential studies.
Set
$\bar{a}_{j,l}=\left\{\begin{array}[]{ll}a_{j,l},&\hbox{if}\ l\in\bar{\mathcal{I}}_{j},\\ 0,&\hbox{otherwise},\end{array}\right.$
and represent $i-1\in\mathcal{D}_{\kappa}$ by the vector $\delta_{\kappa}^{i}$, $i=1,\cdots,\kappa$. Then, $\alpha_{i}(k),u_{i}(k),\bar{a}_{j,l}\in\mathcal{D}_{\kappa}$ have their corresponding vector forms (still using the same symbols) $\alpha_{i}(k),u_{i}(k),\bar{a}_{j,l}\in Col(I_{\kappa})$, respectively. Letting $\alpha(k)=\ltimes_{i=1}^{n}\alpha_{i}(k)$, $u(k)=\ltimes_{i=1}^{n}u_{i}(k)$, we have $\mathcal{C}_{\alpha},\mathcal{C}_{u}(\alpha(k))\subseteq Col(I_{N})$, where $N:=\kappa^{n}$. Then, based on similar calculation to [wangs2022], we have the following result.
Lemma 6.
MAS (4) has the following equivalent algebraic form:
@@ 4.1 Algebraic reformulation of the MAS
where $F\in\mathcal{L}^{N\times N^{2}}$ satisfies $[F]_{:,l}=\ltimes_{j=1}^{n}[F_{j}]_{:,l}$, $l=1,\cdots,N^{2}$, $F_{j}=S_{\kappa}^{n}\ltimes_{l=1}^{n}[I_{\kappa^{l-1}}\otimes(M_{\kappa}\bar{a}_{j,l})]W_{[\kappa,\kappa^{n}]}({\bf{1}}_{\kappa^{j-1}}^{\top}\otimes I_{\kappa}\otimes{\bf{1}}_{\kappa^{n-j}}^{\top})$, $j=1,\cdots,n$, $S_{\kappa}$ and $M_{\kappa}$ respectively represent the structural matrices for $+_{\kappa}$ and ×$\times_{\kappa}$ [Liy2019], ${\bf{1}}_{s}$ represents the $s\times 1$ vector with all entries being $1$, and $W_{[s,t]}$ represents the swap matrix defined as $W_{[s,t]}=[I_{t}\otimes\delta_{s}^{1}~{}\cdots~{}I_{t}\otimes\delta_{s}^{s}]$.
4.2 Criterion of constrained set stabilization
In this part, based on the constrained reachable set, we present a criterion for the constrained set stabilization of MAS (14).
Denote the set of constrained states reachable from $\alpha_{0}$ by $\mathcal{R}(\alpha_{0})$. Then, we have the following result.
Lemma 7.
MAS (14) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, if and only if $I(\Omega(s))\cap\mathcal{R}(\alpha_{0})\neq\emptyset$.

### v2

@@ 3 Optimal wireless control as optimal constrained set stabilization
Given $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$, MAS (4) is constrained $\mathcal{M}$-stabilizable from initial state $\alpha_{0}$, if there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$ such that $\alpha(k)\in\mathcal{C}_{\alpha}$ holds for any $k<K_{0}$, while $\alpha(k)\in\mathcal{M}$ holds for any $k\geq K_{0}$.
@@
A subset $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$ is called a constrained control invariant subset of MAS (4), if for any $\alpha_{0}\in\mathcal{M}$, there exists an admissible input sequence ${\bf{u}}$ such that $\alpha(k)\in\mathcal{M}$ holds for any $k=0,1,\cdots$.
Obviously, the union of any two constrained control invariant subsets of MAS (4) is another constrained control invariant subset. The union of all constrained control invariant subsets contained in $\mathcal{M}$ is called its largest constrained control invariant set, denoted by $I(\mathcal{M})$.
@@
The Lyapunov-like performance (6) is guaranteed for WCS (3) with the state-dependent fading channel (5) and safety constraints (8), if and only if MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, where $\Omega(s)=\{\alpha\in\mathcal{C}_{\alpha}:\bar{\lambda}_{i}(\alpha)\geq s_{i},i=1,\cdots,q\}$, $s=[s_{1}~{}\cdots~{}s_{q}]$ with
$s_{i}=\sup\limits_{y\in\mathbb{R}^{n_{i}},y\neq 0}\frac{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})y}{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})y}.$ (9)
@@
For each subsystem $i$, since $\mathbb{E}[\xi_{i}(l)]=\mathbf{0}$ and $\lambda_{i}(l)$ is independent of $x_{i}(l)$, by (3), it holds
$\mathbb{E}[V_{i}(x_{i}(l+1))|x_{i}(l),\alpha(k),k=\lfloor l/\tau\rfloor]$
$=$ $\mathbb{P}\{\lambda_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}x_{i}^{\top}(l)A_{c,i}^{\top}Q_{i}A_{c,i}x_{i}(l)$
+$+\mathbb{P}\{\lambda_{i}(l)=0|\alpha(k),k=\lfloor l/\tau\rfloor\}x_{i}^{\top}(l)A_{o,i}^{\top}Q_{i}A_{o,i}x_{i}(l)$
@@
Then, condition (6) subject to multiloop WCS dynamics (3) and safety constraints holds, if and only if there exist an admissible input sequence and a positive integer $L_{0}$ such that $\alpha(k)\in\mathcal{C}_{\alpha}$, and for all $x_{i}(l)\neq\mathbf{0}$, it holds
$\mathbb{P}\{\lambda_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}$
$\geq$ $\frac{x_{i}^{\top}(l)(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})x_{i}(l)}{x_{i}^{\top}(l)(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})x_{i}(l)},$ (10)
which are equivalent to the condition that there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$, such that $\alpha(k)\in\mathcal{C}_{\alpha}$, and for all $k\geq K_{0}$, it holds $\bar{\lambda}_{i}(\alpha(k))\geq s_{i}$, $s_{i}$ is given in (9), that is, MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$.
Next, we prove that MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$ if and only if it is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. We just need to prove the necessity. This is proved by a contradiction argument. Since MAS (4) is constrained $\Omega(s)$-stabilizable from $\alpha_{0}$, we can obtain a feasible admissible input sequence ${\bf{u}}^{\prime}$ and a positive integer $K_{0}$, such that for any integer $k<K_{0}$, it holds $\alpha(k;\alpha_{0},{\bf{u}}^{\prime})\in\mathcal{C}_{\alpha}$; for any integer $k\geq K_{0}$, it holds $\alpha(k;\alpha_{0},{\bf{u}}^{\prime})\in\Omega(s)$.
Assuming that MAS (4) is not constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then under input sequence ${\bf{u}}^{\prime}$, there exists an integer $k_{0}\geq K_{0}$ satisfying $\alpha(k_{0})\notin I(\Omega(s))$. Since $I(\Omega(s))$ is a largest constrained control invariant subset, for any $\alpha^{\prime}\in\Omega(s)$ and any admissible input sequence ${\bf{u}}$, if $\alpha^{\prime}\notin I(\Omega(s))$, there exists an integer $k(\alpha^{\prime},{\bf{u}})$ such that $\alpha(k(\alpha^{\prime},{\bf{u}}))\notin\Omega(s)$ from $\alpha^{\prime}$. Then, under input sequence ${\bf{u}}^{\prime}$, it holds $\alpha(k_{0}+k(\alpha(k_{0}),{\bf{u}}^{\prime\prime}))\notin\Omega(s)$ from $\alpha_{0}$, $u^{\prime\prime}(k)=u^{\prime}(k+k_{0})$, which is a contradiction. Thus, MAS (4) is constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$. $\Box$
By (10), $s_{i}$ is the lower bound of the successful transmission probability for system $i$ ensuring the desired Lyapunov-like performance in Definition 1. The value of $s_{i}$ can be obtained by the following linear program:
$\min_{\theta_{i}\geq 0}~{}\theta_{i}$
@@
which can be solved using the toolbox such as YALMIP in MATLAB. In addition, $\Omega(s)$ consists of all constrained states of the MAS satisfying the performance requirements constraint (6) of WCS (3).
For the optimal control problem $P_{1}$, all constraints can be guaranteed if MAS (4) is constrained $I(\Omega(s))$-stabilizable. Since $\mathbb{P}\{\varphi_{i}(l)=1|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}=\sum_{c=0}^{r-1}\bar{\gamma}_{i}(c,\alpha)h_{i}(c)$, ${\bf{u}}^{\ast}$ is a solution to $P_{1}$, if and only if it is a solution to the following infinite-horizon optimal constrained set stabilization problem of MAS (4):
$P_{2}:$ $\min_{{\bf{u}}}\bar{J}_{{\bf{u}}}=\lim_{K\rightarrow\infty}\frac{1}{K}\sum_{k=0}^{K-1}\bar{g}(\alpha(k),u(k))$
[…]
$~{}~{}~{}~{}~{}~{}~{}\,\exists\ K_{0}>0,\alpha(k)\in I(\Omega(s)),\forall k\geq K_{0},$
$~{}~{}~{}~{}~{}~{}~{}\,\alpha(k)\in\mathcal{C}_{\alpha},\forall k<K_{0},$
where $\bar{g}(\alpha(k),u(k))$ is the joint stage cost defined as
$\bar{g}(\alpha(k),\!u(k))\!=\!\tau\!\sum_{i=1}^{q}\!\mu_{i}\!\sum_{c=0}^{r-1}\!\bar{\gamma}_{i}(c,\alpha(k))h_{i}(c)\!+\!\lambda g(\alpha(k),\!u(k)).$
Let the optimal objective values of problems $P_{1}$ and $P_{2}$ be $J^{\ast}=J_{{\bf{u}}^{\ast}}$ and $\bar{J}^{\ast}=\bar{J}_{{\bf{u}}^{\ast}}$, respectively. It is easy to obtain that $J^{\ast}=\bar{J}^{\ast}/\tau$.
4 Infinite-horizon optimal controller design
Note that both the state space and control space of the MAS dynamics (4) are finite. This property allows us to adopt the graphical method to address the optimization problem $P_{2}$. By reformulating the logical-state MAS in a tractable form via the semi-tensor product approach, the constrained optimal state transition graph is firstly constructed to encode the constraints and the objective function of $P_{2}$. Based on this, the feasibility of problem $P_{2}$ is deduced to the existence of cycles in the graph, and the design of optimal input sequences is modeled as a minimum-mean cycle problem. An algorithm is then proposed to construct optimal controllers.
4.1 Feasibility analysis via a graphical method
Due to the highly nonlinear finite-field model, it is difficult to apply classical nonlinear system theory to investigate (4). Recently, the semi-tensor product approach has emerged as an elegant tool to deal with such issues [chengqi2011, cheng2011]. The basic idea is to embed the nonlinear finite-filed dynamics in (4) into a higher dimensional space where its evolution is linear. Specially, representing $i-1\in\mathcal{D}_{\kappa}$ as the vector $\delta_{\kappa}^{i}$, $i=1,\cdots,\kappa$, then $\alpha_{i}(k),u_{i}(k)\in\mathcal{D}_{\kappa}$ have their corresponding vector forms (still using the same symbols) $\alpha_{i}(k),u_{i}(k)\in Col(I_{\kappa})$, respectively. According to similar calculation to [wangs2022], we convert MAS (4) into the following bilinear equivalent algebraic form:
@@ 4.1 Feasibility analysis via a graphical method
where $\alpha(k)=\ltimes_{i=1}^{n}\alpha_{i}(k)\subseteq Col(I_{N})$, $u(k)=\ltimes_{i=1}^{n}u_{i}(k)\subseteq Col(I_{N})$, $N:=\kappa^{n}$, and $F$ is a logical matrix [chengqi2011, cheng2011] characterizing all state-transition information of the MAS. Then, the set of admissible inputs steering MAS (11) from $\delta_{N}^{a}$ to $\delta_{N}^{b}$ in one step can be obtained as
$U_{a,b}=\{\delta_{N}^{l}\in\mathcal{C}_{u}(\delta_{N}^{a}):[F\delta_{N}^{l}]_{b,a}=1\}.$ (12)
In addition, the set of optimal inputs enabling the transition with the lowest cost is
$\bar{U}_{a,b}={\arg\min}_{\delta_{N}^{l}\in U_{a,b}}\bar{g}(\delta_{N}^{a},\delta_{N}^{l}).$ (13)
Figure 3: Constrained optimal state transition graph $G=(\mathcal{V},\mathcal{E},w)$.
To determine the feasibility of optimization problem $P_{2}$, we construct the constrained optimal state transition graph $G=(\mathcal{V},\mathcal{E},w)$ to encode constraints and the objective function. Regard each state $\delta_{N}^{a}\in\mathcal{C}_{\alpha}$ as a vertex on a directed graph, and encode each constrained state transition from $\delta_{N}^{a}$ to $\delta_{N}^{b}$ into edge $(\delta_{N}^{a},\delta_{N}^{b})$. Additionally, assign the weight $w(\delta_{N}^{a},\delta_{N}^{b})=\bar{g}(\delta_{N}^{a},\delta_{N}^{l})$ to the edge $(\delta_{N}^{a},\delta_{N}^{b})\in\mathcal{E}$ to encode the lowest joint cost under the optimal input $\delta_{N}^{l}\in\bar{U}_{a,b}$. See Fig. 3 for an illustration.
For an input sequence ${\bf{u}}$, we say that the state trajectory steered by it converges to a cycle $\mathcal{C}$, if the path $\mathcal{T}=\{\alpha(k):k=0,1,\cdots\}$ in $G$ keeps repeating $\mathcal{C}$ after a finite time. Define the weight and average weight of a path $\mathcal{T}=\{\delta_{N}^{a_{0}},\cdots,\delta_{N}^{a_{l}}\}\subseteq\mathcal{V}$, $l\geq 0$ as $w(\mathcal{T})=\sum_{i=0}^{l-1}w(\delta_{N}^{a_{i}},\delta_{N}^{a_{i+1}})$ and $\bar{w}(\mathcal{T})=w(\mathcal{T})/l$, respectively. Then, we have the following result.
Theorem 5.
MAS (11) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, if and only if $I(\Omega(s))\cap\mathcal{R}(\alpha_{0})\neq\emptyset$, where $\mathcal{R}(\alpha_{0})$ denotes the set of constrained states reachable from $\alpha_{0}$. In addition, if MAS (11) is constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then
  1. (i)
    it holds $C(G[\Phi])\neq\emptyset$, where $\Phi=I(\Omega(s))\cap\mathcal{R}(\alpha_{0})$, $G[\Phi]=(\Phi,\mathcal{E}^{\prime},w)$ is the induced subgraph of $G$ induced by $\Phi$ with =$\mathcal{E}^{\prime}=\{(\delta_{N}^{a},\delta_{N}^{b})\in\mathcal{E}:\delta_{N}^{a},\delta_{N}^{b}\in\Phi\}$, and $C(\cdot)$ represents the set of all cycles in the graph;
  2. (ii)
    solutions to the infinite-horizon optimal control problem $P_{2}$ exist.